OpenStudy (jpsmarinho):

The derivative of arccossec(x) is:

OpenStudy (jpsmarinho):

is that correct?

OpenStudy (anonymous):

\[\frac{d}{dx}f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

derivative of cosecant is \(-\csc(x)\cot(x)\) if i remember correctly, so your answer is \[-\frac{1}{\csc(\csc^{-1}(x))\cot(\csc^{-1}(x))}\] but you can clean up the denominator and leave no trace of a trig function

OpenStudy (anonymous):

now that i look at your answer, it is right, but you need to replace \(y\) by \(\csc^{-1}(x)\) to complete it

OpenStudy (jpsmarinho):

But, y is equals to arccossec(x) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and further \(\csc(\csc^{-1}(x))=x\) and also \(\cot(\csc^{-1}(x))=\sqrt{x^2-1}\)

OpenStudy (jpsmarinho):

Aah when you told csc^-1(x) , I think that you wrote 1/csc(x). Thanks!

OpenStudy (anonymous):

yeah those mean inverse functions

OpenStudy (anonymous):

so your "final answer" is \(-\frac{1}{x\sqrt{x^2-1}}\)

OpenStudy (jpsmarinho):

Wait, @satellite73 , please. How I discover that? cot(arccsc(x)) is equals to (x^2 - 1)^(1/2)? It's something about the identities sec(x) squared = tan(x) squared + 1?